OpenStudy (anonymous):

Calculus? Anti derivatives? Please help? verify the statement by showing that the derivative of the right side equals the integrand of the left side. Integral (4x^3 - 1/x^2) dx = x^4 + 1/x + C

OpenStudy (anonymous):

Take derivative of the right side, let say it is = "something" , you can see that that"something" is the integrand of the left side The verification is done

OpenStudy (anonymous):

The derivative on the right side: (x^4 + 1/x + C)' = (x^4)' + (1/x)' + C' = 4x^3 - 1/x^2.

OpenStudy (anonymous):

ahhh ok

OpenStudy (anonymous):

Thank you!